OpenStudy (anonymous):

rewrite the equation 2x^2-3=-6x-2 to be in the form ax^2+bx+c=0

OpenStudy (anonymous):

@satellite73 can you please help me?

OpenStudy (anonymous):

I will give a medal to whoever can help me! :) the options are A)2x^2-6x-1=0 B) 2x^2+6x-5=0 c) 2x^2+6x-1=0 D) 2x^2-6x-5=0

OpenStudy (justuu):

I think its C

OpenStudy (anonymous):

yeah thanks!

OpenStudy (justuu):

No problem

OpenStudy (ajspeller):